# Varying coefficients correlated velocity models in complex landscapes with boundaries applied to narwhal responses to noise exposure

### v3

@@ 1 Introduction
There is a large variety of sources of anthropogenic underwater noise in the Arctic region, including sonars, ice breakers, vessel traffic, drilling and seismic airguns used for oil and gas exploration (Halliday, Pine and Insley, 2020). A better understanding of the impact of these noises on marine life is critical for conservation policies. Marine mammals, whose vital functions highly depend on sound perception for communication and orientation, are particularly vulnerable. Besides causing physical harm, anthropogenic noise can also disturb their behavior and interrupt their natural foraging habits. However, there is still no clear criteria for measuring behavioral disturbances due to the multiplicity of contextual variables that can influence a behavioral reaction, and the variety of these reactions (Southall et al., 2008, 2019). Comprehensive behavioral studies are thus essential.
Several surveys of behavioral responses have been conducted on beaked whales, which have been regularly involved in stranding events following sonar exercises (Tyack et al., 2011; Cioffi et al., 2022). Studies have also been conducted on belugas (Martin et al., 2023), sperm whales (Madsen et al., 2006), blue whales (Friedlaender et al., 2016), and narwhals (Heide-Jørgensen et al., 2021; Tervo et al., 2021, 2023). The most critical behavioral responses - those potentially altering a population’s capacity to survive, reproduce or forage - include changes in movement speed and direction, avoidance reactions as well as modified dive profiles or vocalizations (Southall et al., 2008). Studies have shown that narwhals can exhibit some of these critical behavioral responses to sound exposure, such as avoidance reactions or changes of direction to move towards the shore (Heide-Jørgensen et al., 2021). A significant decrease in their buzzing rate has also been assessed as far as 40 km away from the sound source (Tervo et al., 2021). These results are based on controlled exposure experiments conducted in 2017 and 2018 in a pristine area in South-East Greenland, still largely unaffected by noise pollution, and which is home to a declining population of narwhals (Garde et al., 2022). Our study is based on these data collected in 2018.
@@
We specifically focus on continuous-time models defined as a solution of a stochastic differential equation (SDE) with time-varying coefficients. One widely known model in this context is the continuous-time correlated random walk, also known as the integrated Ornstein-Uhlenbeck or correlated velocity model (CVM) (Johnson et al., 2008). It can be viewed as a special case of velocity potential models, where the potential is a second degree polynomial (Preisler, Ager and Wisdom, 2013). Varieties of CVMs have been applied to birds (Janaswamy, Loring and McLaren, 2018), marine algae (Gurarie, Grünbaum and Nishizaki, 2011), ants (Russell et al., 2018), seals (Johnson et al., 2008; Albertsen, 2018), bowhead whales (Gurarie, Grünbaum and Nishizaki, 2011) and sea lion motion (Hanks, Johnson and Hooten, 2017).
@@
In our case study, the narwhals move in a restricted domain that consists in a complex system of fjords in Scoresby Sound (South-East Greenland), the largest fjord system in the world, known to be the summer residence for an isolated population of narwhals. Thus, the modeling process must include shoreline boundaries through movement constraints. Most existing CVMs for animal movement do not account for such constraints (Johnson et al., 2008; Michelot et al., 2021; Gurarie et al., 2017). Examples models including landscape boundaries are constrained SDEs (Hanks, Johnson and Hooten, 2017) in which a reflected version of the CVM is considered to study sea lion telemetry data, and SDEs with drift described by a potential function (Russell et al., 2018) where a repulsive potential function constrains ants movement within a box. (Brillinger, 2003) also illustrates different methods of simulation for constrained animal movement, and emphasizes that the way to model motion near the boundary of the domain should be species-specific and account for the particular behaviors and spatial use of the studied animal. Narwhals have a proclivity to rotate to avoid or align with the shoreline without reaching the boundary. A reflected SDE as in (Hanks, Johnson and Hooten, 2017) does not allow this type of behavior. In addition, a simple exponential potential function is suitable only with basic boundaries, such as the rectangular box in (Russell et al., 2018), but would break down for a complex boundary such as Scoresby Sound fjords. As a consequence, there is a need to develop new models adapted to such spatial constraints and behaviors. Reflected SDEs have certain limitations. On the one hand, it is not a modelling of shoreline avoidance behaviour. The model only uses a projection operator to ensure that the process remains within the domain. On the other hand, the transition density of these models is not known, which complicates the statistical inference of the parameters. We therefore propose to work with the second class of model, SDEs whose drift constraints movement within the system of fjords.
@@
The novelty of this paper consists in extending the rotational CVM defined in (Gurarie et al., 2017) with a rotation parameter expressed as a smooth function of the distance to the boundary and the angle between the animal’s heading and the boundary normal vector. This makes the drift in the velocity equation dependent on the location process and the boundary of the domain, and allows the velocity to rotate as the narwhal approaches the shore. We show that the rotational parameter can be estimated as a function of the two covariates in the framework of the R package smoothSDE (Michelot et al., 2021). To make this possible, we derive explicit formulas for the transition density of the location and velocity process and formulate a linear Gaussian state-space model from the discretisation of these formulas. Maximum likelihood estimation based on the Kalman filter is then performed as for any other model in smoothSDE. The estimation procedure exploits the capabilities of the R package TMB to get approximate likelihood using Laplace’s approximation (Kristensen et al., 2016; Albertsen et al., 2015). It is considerably less computationally intensive than an estimation procedure based on Markov Chain Monte Carlo posterior samples as done in (Hanks, Johnson and Hooten, 2017; Russell et al., 2018).
@@
In this general framework, sound exposure is introduced in the model through smooth functions of a covariate defined as the inverse of the distance to the sound source. This choice is motivated by (Heide-Jørgensen et al., 2021; Tervo et al., 2021, 2023). We estimate the effect of the exposure variables on both the velocity and the persistence parameters of the diffusion process. A response model is compared to a baseline model for the narwhal movement before exposure to the disturbance, that is, under normal conditions (Michelot et al., 2022). The effect of sound exposure is assessed as a deviation in the response model parameters when compared to the baseline model parameters.
@@
  • Derivation of explicit formulas for the transition density of the rotational CVM defined in (Gurarie et al., 2017), and addition of this model in the framework of the R package smoothSDE, to enable the use of smooth parameters depending on external covariates, and estimation from noisy observations irregularly spaced in time.
@@ 2.1 Description of the controlled exposure experiments
The dataset analysed in this paper has been subject of several studies, focusing mainly on vocalizations and avoidance reactions (Heide-Jørgensen et al., 2021; Tervo et al., 2021, 2023). Here we briefly describe the data, for more details see (Heide-Jørgensen et al., 2021). Six male narwhals were equipped with FastLoc GPS receivers in August 2018 in Scoresby Sound in South-East Greenland by biologists from the Greenland Institute of Natural Ressources, with the help of local Inuit hunters. An offshore patrol vessel military ship was sailed to shoot airguns underwater between August 25 and September 1. It was equipped with two airguns at 6m depth and moved at a speed of 4.5 knots. The guns were fired synchronously every $80$ s, while the GPS navigation system recorded the location of every shot. The data include latitude and longitude of the narwhal positions, distance relative to the ship in metres, GPS time, and distance to the shore in metres. GPS positions are known only at times when the narwhals are at the surface. The median time step between two GPS measurements in the data is about $5$ minutes and only $0.3\%$ of the time steps reach more than two hours, with a maximum at more than $4$ hours. While the statistical analysis in (Heide-Jørgensen et al., 2021) relied on augmented data with positions linearly interpolated at each second, here we only consider the actual GPS measurements with irregular time steps.
For each narwhal, the entire track is split into two periods: a period before exposure defined as the period before the narwhal gets in line of sight with the ship for the first time, and the period of exposure starting after exposure onset. We discard the GPS measurements from the first 12 hours after tagging, to avoid any tagging effects on the behavior. Measurements that result in a velocity higher than $20$ km/h were also discarded (only 2 data points for the same narwhal track). Overall, $4815$ GPS positions are kept for the analysis. The splitting between data before and after exposure results in $1558$ measurements before exposure and $3257$ measurements after exposure.
@@ 2.2 Notation and definition of covariates
The distance to the ship $D^{ship}$ (km) is defined as the distance between observed GPS locations of the ship and the narwhal. The values are comprised between $2.68$ and $63.8$ km. Exposure to the ship for narwhal $i$ at time $t_{ij}$, denoted $E^{ship}_{ij}$, is defined as the inverse of the distance to the ship (in km) (Heide-Jørgensen et al., 2021). The covariate $E^{ship}$ is meant to be a proxy for sound exposure levels received by the narwhals. Reasonably, the closer the narwhal is to the ship, the louder is the sound, and thus the exposure is larger. $D^{ship}$ is only defined when the narwhal is in line of sight with the ship and exposures are set to 00 when this is not this case. This implies that in the statistical model, the ship noise is only allowed to affect the narwhal when the ship is in line of sight, though it is likely that narwhals can perceive the disturbance even when they are not in line of sight with the sound source. This provides a conservative estimate of the effect of the noise exposure. Levels of the covariates $I^{shore}$ and $E^{ship}$ for each narwhal are displayed in Figure 2.
@@ 3.1 Standard Rotational Advective correlated velocity model (RACVM)
where $V(t)$ is the horizontal two-dimensional velocity at time $t$, $X(t)$ is the two-dimensional horizontal position, typically in longitude and latitude or in UTM coordinates, $W(t)$ is a two-dimensional brownian motion. Parameter $\tau$ is an autocorrelation time scale, $\nu$ controls the norm of the velocity and drives the random variability, $\mu=\begin{pmatrix}\mu_{1}&&\mu_{2}\end{pmatrix}^{\top}$ is the long term velocity, and $\omega$ is an angular velocity that controls how fast the velocity vector rotates. The case $\omega=0$ corresponds to a standard CVM. In most applications, this standard model is satisfactory, but in case of tortuous movement, a non-zero value of $\omega$ is sometimes needed (Gurarie et al., 2017; Alt, Hoffmann and Levin, 1990; Albertsen, 2018). Moreover, whether or not to include a non-zero mean velocity parameter $\mu$ depends on the specific context of the study. Typically, it makes sense to have $\mu\neq(0,0)$ when examining migratory patterns or avoidance reactions from a fixed sound source. We refer to (2) as ”Rotational Correlated Velocity Model” (RCVM) when $\mu=(0,0)$, and ”Rotational Advective Correlated Velocity Model” (RACVM) when $\mu\neq(0,0)$.
In this section, we derive the explicit transition density for the process $U=\begin{pmatrix}X&&V\end{pmatrix}^{\top}$. Though this model is considered in several papers, we found no mention of the closed form formulas derived here. In Gurarie et al. (2017), only the process $V$ is exhaustively studied, and the classical distributional results about this process are used for estimation and simulation purposes. A more general formulation is considered by Albertsen (2018) in which the diffusion matrix is lower triangular with positive diagonal elements and two distinct autocorrelation parameters are considered in each direction allowing for anisotropic movement. In this case, Albertsen (2018) derives the Gaussian transition density of the velocity process, where the covariance matrix is expressed using a Kronecker sum. However, the formulas for the process $X$ are not provided, and the Euler scheme is employed to approximate the transition densities of $X$. This approach is generally sufficient when the animal’s positions are recorded at high frequency. Yet, for marine mammals such as narwhals that dive to great depths, GPS measurements are usually taken at irregular intervals, typically of several minutes, which can render the Euler scheme unsuitable. Additionally, incorporating covariates into the parameters of the SDE often requires approximating the transition density by assuming the covariates remain constant during each time step (Michelot et al., 2021). This results in two layers of approximation, possibly leading to unreliable estimation.
@@
The formulas derived in (Johnson et al., 2008) are a corollary of Proposition 3.1, obtained with $\omega=0$. The proof of this proposition is detailed in Appendix.
@@ 3.2 Constrained rotational correlated velocity model
The $\psi_{k},k=1,\ldots,L$ can be constructed in several ways based on thin plate regression splines, or on the tensor product of any univariate spline basis. The coefficients $\omega_{k}$ control the contribution of each basis function to the final shape of the smooth function (Wood, 2017). Figure 3(a) shows function (8) in the top, along with a smooth function $f_{\omega}$ obtained as a combination of $L=9$ cubic spline basis functions approximating (8) in the bottom plots.
@@ 4.1 Mixed effect CRCVM model for shore influence and sound exposure
The standard deviation $\sigma_{obs}$ represents GPS measurement errors variability. The assumption of i.i.d. Gaussian measurement errors is a simplified model of actual GPS measurement errors. In practice, the accuracy of GPS positions is influenced by the number of satellites processing the GPS signal—more satellites generally result in more accurate positions. Furthermore, the error can differ between the $x$ and $y$ directions. For a comprehensive study on Fastloc-GPS measurement errors, we refer to (Wensveen, Thomas and Miller, 2015). It is often more appropriate to use a distribution with heavier tails, such as the Student’s $t$-distribution, to model measurement error. However, this introduces additional complexity to the inference process, as a standard linear Kalman filter cannot be directly applied in such cases. Since the Student’s $t$-distribution approaches a Gaussian distribution as its degrees of freedom increase, the Gaussian distribution is considered a reasonable trade-off between simplicity and suitability.
@@
Since the two covariates are not on the same scale (one is in $rad$ while the other is in $km^{-1}$), tensor splines are more appropriate than thin plate regression splines. The functions $\psi_{k}$ in equation (12) are the bivariate basis functions constructed from univariate cubic spline basis functions. We refer to (Wood, 2017) section $4.1.8$ for the construction of tensor splines basis functions. The splines degree of freedom is $L=q_{E}\times q_{\Theta}$ and the number of knots in the splines are $q_{E}-1$ and $q_{\Theta}-1$ for the covariates $I_{shore}$ and $\Theta$, respectively.

### v5

@@ 1 Introduction
There is a large variety of sources of anthropogenic underwater noise in the Arctic region, including sonars, ice breakers, vessel traffic, drilling and seismic airguns used for oil and gas exploration (halliday_underwater_2020). A better understanding of the impact of these noises on marine life is critical for conservation policies. Marine mammals, whose vital functions highly depend on sound perception for communication and orientation, are particularly vulnerable. Besides causing physical harm, anthropogenic noise can also disturb their behavior and interrupt their natural foraging habits. However, there is still no clear criteria for measuring behavioral disturbances due to the multiplicity of contextual variables that can influence a behavioral reaction, and the variety of these reactions (southall_marine_2008; southall_marine_2019). Comprehensive behavioral studies are thus essential.
Several surveys of behavioral responses have been conducted on beaked whales, which have been regularly involved in stranding events following sonar exercises (tyack_beaked_2011; cioffi_trade-offs_2022). Studies have also been conducted on belugas (martin_exposure_2023), sperm whales (madsen_quantitative_2006), blue whales (friedlaender_preymediated_2016), and narwhals (heide-jorgensen_behavioral_2021; tervo_narwhals_2021; tervo_stuck_2023). The most critical behavioral responses - those potentially altering a population’s capacity to survive, reproduce or forage - include changes in movement speed and direction, avoidance reactions as well as modified dive profiles or vocalizations (southall_marine_2008). Studies have shown that narwhals can exhibit some of these critical behavioral responses to sound exposure, such as avoidance reactions or changes of direction to move towards the shore (heide-jorgensen_behavioral_2021). A significant decrease in their buzzing rate has also been assessed as far as 40 km away from the sound source (tervo_narwhals_2021). These results are based on controlled exposure experiments conducted in 2017 and 2018 in a pristine area in South-East Greenland, still largely unaffected by noise pollution, and which is home to a declining population of narwhals (garde_biological_2022). Our study is based on these data collected in 2018.
@@
We specifically focus on continuous-time models defined as a solution of a stochastic differential equation (SDE) with time-varying coefficients. One widely known model in this context is the continuous-time correlated random walk, also known as the integrated Ornstein-Uhlenbeck or correlated velocity model (CVM) (johnson_continuous_2008). It can be viewed as a special case of velocity potential models, where the potential is a second degree polynomial (preisler_analyzing_2013). Varieties of CVMs have been applied to birds (janaswamy_state_2018), marine algae (gurarie_estimating_2011), ants (russell_spatially_2018), seals (johnson_continuous_2008; albertsen_generalizing_2018), bowhead whales (gurarie_estimating_2011) and sea lion motion (hanks_reflected_2017).
@@
In our case study, the narwhals move in a restricted domain that consists in a complex system of fjords in Scoresby Sound (South-East Greenland), the largest fjord system in the world, known to be the summer residence for an isolated population of narwhals. Thus, the modeling process must include shoreline boundaries through movement constraints. Most existing CVMs for animal movement do not account for such constraints (johnson_continuous_2008; michelot_varying-coefficient_2021; gurarie_correlated_2017). Example models including landscape boundaries are constrained SDEs (hanks_reflected_2017) in which a reflected version of the CVM is considered to study sea lion telemetry data, and SDEs with drift described by a potential function (russell_spatially_2018) where a repulsive potential function constrains ants movement within a box. (brillinger_simulating_2003) also illustrates different methods of simulation for constrained animal movement, and emphasizes that the way to model motion near the boundary of the domain should be species-specific and account for the particular behaviors and spatial use of the studied animal. Narwhals have a proclivity to rotate to avoid or align with the shoreline without reaching the boundary. A reflected SDE as in (hanks_reflected_2017) does not allow this type of behavior. In addition, a simple exponential potential function is suitable only with basic boundaries, such as the rectangular box in (russell_spatially_2018), but would break down for a complex boundary such as Scoresby Sound fjords. As a consequence, there is a need to develop new models adapted to such spatial constraints and behaviors. Reflected SDEs have certain limitations. On the one hand, it is not a modelling of shoreline avoidance behaviour. The model only uses a projection operator to ensure that the process remains within the domain. On the other hand, the transition density of these models is not known, which complicates the statistical inference of the parameters. We therefore propose to work with the second class of model, SDEs whose drift constraints movement within the system of fjords.
@@
The novelty of this paper consists in extending the rotational CVM defined in (gurarie_correlated_2017) with a rotation parameter expressed as a smooth function of the distance to the boundary and the angle between the animal’s heading and the boundary normal vector. This makes the drift in the velocity equation dependent on the location process and the boundary of the domain, and allows the velocity to rotate as the narwhal approaches the shore. We show that the rotational parameter can be estimated as a function of the two covariates in the framework of the R package smoothSDE (michelot_varying-coefficient_2021). To make this possible, we derive explicit formulas for the transition density of the location and velocity process and formulate a linear Gaussian state-space model from the discretisation of these formulas. Maximum likelihood estimation based on the Kalman filter is then performed as for any other model in smoothSDE. The estimation procedure exploits the capabilities of the R package TMB to get approximate likelihood using Laplace’s approximation (kristensen_tmb_2016; albertsen_fast_2015). It is considerably less computationally intensive than an estimation procedure based on Markov Chain Monte Carlo posterior samples as done in (hanks_reflected_2017; russell_spatially_2018).
@@
In this general framework, sound exposure is introduced in the model through smooth functions of a covariate defined as the inverse of the distance to the sound source. This choice is motivated by (heide-jorgensen_behavioral_2021; tervo_narwhals_2021; tervo_stuck_2023). We estimate the effect of the exposure variables on both the velocity and the persistence parameters of the diffusion process. A response model is compared to a baseline model for the narwhal movement before exposure to the disturbance, that is, under normal conditions (michelot_continuous-time_2022). The effect of sound exposure is assessed as a deviation in the response model parameters when compared to the baseline model parameters.
@@
  • Derivation of explicit formulas for the transition density of the rotational CVM defined in (gurarie_correlated_2017), and addition of this model in the framework of the R package smoothSDE, to enable the use of smooth parameters depending on external covariates, and estimation from noisy observations irregularly spaced in time.
@@ 2.1 Description of the controlled exposure experiments
The dataset analysed in this paper has been subject of several studies, focusing mainly on vocalizations and avoidance reactions (heide-jorgensen_behavioral_2021; tervo_narwhals_2021; tervo_stuck_2023). Here we briefly describe the data, for more details see (heide-jorgensen_behavioral_2021). Six male narwhals were equipped with FastLoc GPS receivers in August 2018 in Scoresby Sound in South-East Greenland by biologists from the Greenland Institute of Natural Ressources, with the help of local Inuit hunters. An offshore patrol vessel military ship was sailed to shoot airguns underwater between August 25 and September 1. It was equipped with two airguns at 6m depth and moved at a speed of 4.5 knots. The guns were fired synchronously every $80$ s, while the GPS navigation system recorded the location of every shot. The data include latitude and longitude of the narwhal positions, distance relative to the ship in metres, GPS time, and distance to the shore in metres. GPS positions are known only at times when the narwhals are at the surface. The median time step between two GPS measurements in the data is about $5$ minutes and only $0.3\%$ of the time steps reach more than two hours, with a maximum at more than $4$ hours. While the statistical analysis in (heide-jorgensen_behavioral_2021) relied on augmented data with positions linearly interpolated at each second, here we only consider the actual GPS measurements with irregular time steps.
For each narwhal, the entire track is split into two periods: a period before exposure defined as the period before the narwhal gets in line of sight with the ship for the first time, and the period of exposure starting after exposure onset. We discard the GPS measurements from the first 12 hours after tagging, to avoid any tagging effects on the behavior. Additionally, a velocity filter was applied to retain only positions with an empirical velocity below 20 km/h. This filter removed two data points from the dataset. Overall, $4815$ GPS positions are kept for the analysis. The splitting between data before and after exposure results in $1558$ measurements before exposure and $3257$ measurements after exposure.
@@ 2.2 Notation and definition of covariates
The distance to the ship $D^{ship}$ (km) is defined as the distance between observed GPS locations of the ship and the narwhal. The values are comprised between $2.68$ and $63.8$ km. Exposure to the ship for narwhal $i$ at time $t_{ij}$, denoted $E^{ship}_{ij}$, is defined as the inverse of the distance to the ship (in km) (heide-jorgensen_behavioral_2021). The covariate $E^{ship}$ is meant to be a proxy for sound exposure levels received by the narwhals. Reasonably, the closer the narwhal is to the ship, the louder is the sound, and thus the exposure is larger. $D^{ship}$ is only defined when the narwhal is in line of sight with the ship and exposures are set to 00 when this is not this case. This implies that in the statistical model, the ship noise is only allowed to affect the narwhal when the ship is in line of sight, though it is likely that narwhals can perceive the disturbance even when they are not in line of sight with the sound source. This provides a conservative estimate of the effect of the noise exposure. Levels of the covariates $I^{shore}$ and $E^{ship}$ for each narwhal are displayed in Figure 2.
@@ 3.1 Standard Rotational Advective correlated velocity model (RACVM)
where $V(t)$ is the horizontal two-dimensional velocity at time $t$, $X(t)$ is the two-dimensional horizontal position, typically in longitude and latitude or in UTM coordinates, $W(t)$ is a two-dimensional brownian motion. Parameter $\tau$ is an autocorrelation time scale, $\nu$ controls the norm of the velocity and drives the random variability, $\mu=\begin{pmatrix}\mu_{1}&&\mu_{2}\end{pmatrix}^{\top}$ is the long term velocity, and $\omega$ is an angular velocity that controls how fast the velocity vector rotates. The case $\omega=0$ corresponds to a standard CVM. In most applications, this standard model is satisfactory, but in case of tortuous movement, a non-zero value of $\omega$ is sometimes needed (gurarie_correlated_2017; alt_correlation_1990; albertsen_generalizing_2018). Moreover, whether or not to include a non-zero mean velocity parameter $\mu$ depends on the specific context of the study. Typically, it makes sense to have $\mu\neq(0,0)$ when examining migratory patterns or avoidance reactions from a fixed sound source. We refer to (2) as ”Rotational Correlated Velocity Model” (RCVM) when $\mu=(0,0)$, and ”Rotational Advective Correlated Velocity Model” (RACVM) when $\mu\neq(0,0)$.
In this section, we derive the explicit transition density for the process $U=\begin{pmatrix}X&&V\end{pmatrix}^{\top}$. Though this model is considered in several papers, we found no mention of the closed form formulas derived here. In gurarie_correlated_2017, only the process $V$ is exhaustively studied, and the classical distributional results about this process are used for estimation and simulation purposes. A more general formulation is considered by albertsen_generalizing_2018 in which the diffusion matrix is lower triangular with positive diagonal elements and two distinct autocorrelation parameters are considered in each direction allowing for anisotropic movement. In this case, albertsen_generalizing_2018 derives the Gaussian transition density of the velocity process, where the covariance matrix is expressed using a Kronecker sum. However, the formulas for the process $X$ are not provided, and the Euler scheme is employed to approximate the transition densities of $X$. This approach is generally sufficient when the animal’s positions are recorded at high frequency. Yet, for marine mammals such as narwhals that dive to great depths, GPS measurements are usually taken at irregular intervals, typically of several minutes, which can render the Euler scheme unsuitable. Additionally, incorporating covariates into the parameters of the SDE often requires approximating the transition density by assuming the covariates remain constant during each time step (michelot_varying-coefficient_2021). This results in two layers of approximation, possibly leading to unreliable estimation.
@@
The formulas derived in (johnson_continuous_2008) are a corollary of Proposition 3.1, obtained with $\omega=0$. The proof of this proposition is detailed in Appendix.
@@ 3.2 Constrained rotational correlated velocity model
The $\psi_{k},k=1,\ldots,L$ can be constructed in several ways based on thin plate regression splines, or on the tensor product of any univariate spline basis. The coefficients $\omega_{k}$ control the contribution of each basis function to the final shape of the smooth function (wood_generalized_2017). Figure 3(a) shows function (8) in the top, along with a smooth function $f_{\omega}$ obtained as a combination of $L=9$ cubic spline basis functions approximating (8) in the bottom plots.
@@ 4.1 Mixed effect CRCVM model for shore influence and sound exposure
The standard deviation $\sigma_{obs}$ represents GPS measurement errors variability. The assumption of i.i.d. Gaussian measurement errors is a simplified model of actual GPS measurement errors. In practice, the accuracy of GPS positions is influenced by the number of satellites processing the GPS signal—more satellites generally result in more accurate positions. Furthermore, the error can differ between the $x$ and $y$ directions. For a comprehensive study on Fastloc-GPS measurement errors, we refer to (wensveen_path_2015). It is often more appropriate to use a distribution with heavier tails, such as the Student’s $t$-distribution, to model measurement error. However, this introduces additional complexity to the inference process, as a standard linear Kalman filter cannot be directly applied in such cases. Since the Student’s $t$-distribution approaches a Gaussian distribution as its degrees of freedom increase, the Gaussian distribution is considered a reasonable trade-off between simplicity and suitability.
@@
Since the two covariates are not on the same scale (one is in $rad$ while the other is in $km^{-1}$), tensor splines are more appropriate than thin plate regression splines. The functions $\psi_{k}$ in equation (12) are the bivariate basis functions constructed from univariate cubic spline basis functions. We refer to (wood_generalized_2017) section $4.1.8$ for the construction of tensor splines basis functions. The splines degree of freedom is $L=q_{E}\times q_{\Theta}$ and the number of knots in the splines are $q_{E}-1$ and $q_{\Theta}-1$ for the covariates $I_{shore}$ and $\Theta$, respectively.